OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6. Find (fog)(4). Please explain solution

OpenStudy (darkprince14):

(fog)(x) = f(g(x)) which implies that you have to replace all the X in F(X) by g(x) hope this helps :)

OpenStudy (anonymous):

Sorry no, I need to see how to start the solution.

OpenStudy (anonymous):

\[fog(x)=f(g(x))=3(g(x))-7=3(-2x-6)-7\] Simplify and put x=4

OpenStudy (anonymous):

so, 3(-8-6)-7?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Then 3(-2)-7= -6-7=-1?

OpenStudy (anonymous):

Nope, what is -8-6?

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

Yes, try the calculation again now

OpenStudy (anonymous):

Ok 3(-14)-7=-42-7=-35?

OpenStudy (anonymous):

Nope, you did the same mistake again, Remember it's -42-7 and NOT -42-(-7)

OpenStudy (anonymous):

-49

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome